# Accelerating Goal-Conditioned Reinforcement Learning Algorithms and Research

### v3

@@ Abstract
Self-supervision has the potential to transform reinforcement learning (RL), paralleling the breakthroughs it has enabled in other areas of machine learning. While self-supervised learning in other domains aims to find patterns in a fixed dataset, self-supervised goal-conditioneds reinforcement learning (GCRL) agents discover new behaviors by learning from the goals achieved during unstructured interaction with the environment. However, these methods have failed to see similar success, both due to a lack of data from slow environment simulations as well as a lack of stable algorithms. We take a step toward addressing both of these issues by releasing a high-performance codebase and benchmark (JaxGCRL) for self-supervised GCRL, enabling researchers to train agents for millions of environment steps in minutes on a single GPU. By utilizing GPU-accelerated replay buffers, environments, and a stable contrastive RL algorithm, we reduce training time by up to $\times2$. Additionally, we assess key design choices in contrastive RL, identifying those that most effectively stabilize and enhance training performance. With this approach, we provide a foundation for future research in self-supervised GCRL, enabling researchers to quickly iterate on new ideas and evaluate them in diverse and challenging environments. Code: https://github.com/MichalBortkiewicz/JaxGCRL.
@@ 1 Introduction
Self-supervised learning has significantly influenced machine learning over the last decade, transforming how research is done in domains such as natural language processing and computer vision (Chen et al., 2020b, ; Dosovitskiy et al.,, 2020; Vaswani et al.,, 2017). In the context of reinforcement learning (RL), most self-supervised prior methods apply the same recipe that has been successful in other domains: learning representations or models from a large, fixed dataset (Hoffmann et al.,, 2022; Sardana et al.,, 2024; Muennighoff et al.,, 2023). However, the RL setting also enables a fundamentally different type of self-supervised learning: rather than learning from a fixed dataset (as done in NLP and computer vision), a self-supervised reinforcement learner can collect its own dataset. Thus, rather than learning a representation of a dataset, the self-supervised reinforcement learner acquires a representation of an environment or of behaviors and optimal policies therein. Self-supervised reinforcement learners may address many of the challenges that stymie today’s foundation’s models: reasoning about the consequences of actions (Rajani et al.,, 2019; Kwon et al.,, 2023) (i.e., counterfactuals (Bhargava and Ng,, 2022; Jin et al.,, 2023)) and long horizon planning (Bhargava and Ng,, 2022; Du et al.,, 2023; Guan et al.,, 2023).
@@
In this paper, we study self-supervised RL in an online setting: an agent interacts in an environment without a reward to learn representations, which are later used to quickly solve downstream tasks. We focus on goal-conditioned reinforcement learning (GCRL) algorithms, which aim to use these unsupervised interactions to learn policies for achieving various goals – an essential capability for multipurpose robots. Prior work has proposed several algorithms for self-supervised RL (Eysenbach et al.,, 2022; Zheng et al.,, 2024; Myers et al.,, 2023; Myers et al., 2024a, ), including algorithms that focus on learning goals (Erraqabi et al.,, 2022). However, these methods were limited by small datasets and infrequent online interactions with the environment, which has prevented them from exploring the potential emergent properties of self-supervised reinforcement learning on large-scale data.
The main goal of this paper is to introduce JaxGCRL: an extremely fast GPU-accelerated codebase and benchmark for effective self-supervised GCRL research. For instance, an experiment with $10$ million environment steps lasts only around 10 minutes on a single GPU, which is up to 22$\times$ faster than in the original contrastive RL codebase (Eysenbach et al.,, 2022) (Fig. 1). This speed allows researchers to “interactively” debug their algorithms, changing pieces and getting results in near real-time for multiple random seeds on a single GPU without the hustle of distributed training. Consequently, JaxGCRL eliminates the barriers to entry to state-of-the-art GCRL research, making it more accessible to under-resourced institutions.
To achieve this training and performance improvement, we combine insights from self-supervised RL with recent advances in GPU-accelerated simulation. The first key ingredient is recent work on GPU-accelerated simulators, both for physics (Freeman et al.,, 2021; Thibault et al.,, 2024; Liang et al.,, 2018) and other tasks (Matthews et al.,, 2024; Dalton et al.,, 2020; Fischer et al.,, 2009; Bonnet et al.,, 2024; Rutherford et al.,, 2024) that enable users to collect data up to 1000$\times$ faster (Freeman et al.,, 2021) than prior methods based on CPUs. The second key ingredient is a highly stable algorithm build upon recent work on contrastive RL (CRL) (Eysenbach et al.,, 2022; Zheng et al.,, 2024), which uses temporal contrastive learning to learn a value function. The third key ingredient is a suite of tasks for evaluating self-supervised RL agents which is not only blazing fast but also stress tests the exploration and long-horizon reasoning capabilities of RL policies.
@@ 2.1 Goal-Conditioned Reinforcement Learning
GCRL is a special case of the general multi-task RL setting, in which the potential tasks are defined by goal states that the agent is trying to reach in an environment (Kaelbling,, 1993; Ghosh et al.,, 2019; Nair et al.,, 2018). Achieving any goal is appealing for generalist agents, as it allows for diverse behaviors without needing specific reward functions for each task (each state defines a task when seen as a goal) (Schaul et al.,, 2015). GCRL techniques have seen success in domains such as robotic manipulation (Andrychowicz et al.,, 2017; Eysenbach et al.,, 2021; Ghosh et al.,, 2021; Ding et al.,, 2022; Walke et al.,, 2023) and navigation (Shah et al.,, 2023; Levine and Shah,, 2023; Manderson et al.,, 2020; Hoang et al.,, 2021). Recent work has shown that representations of goals can be imbued with additional structure to enable capabilities such as language grounding (Myers et al.,, 2023; Ma et al.,, 2023), compositionality (Liu et al.,, 2023; Myers et al., 2024b, ; Wang et al.,, 2023), and planning (Park et al.,, 2023; Eysenbach et al.,, 2024). We show how GCRL techniques based on contrastive learning (Eysenbach et al.,, 2022) can be scaled with GPU acceleration to enable fast and stable training.
@@ 2.2 Accelerating Deep Reinforcement Learning
Deep RL has only recently become practical for many tasks, in part due to improvements in hardware support for these algorithms. Distributed training has enabled RL algorithms to scale across hundreds of GPUs (Mnih et al.,, 2016; Espeholt et al.,, 2018; 2020; Hoffman et al.,, 2022). To resolve the bottleneck of environment interaction with CPU-bound environments, various GPU-accelerated environments have been proposed (Matthews et al.,, 2024; Freeman et al.,, 2021; Bonnet et al.,, 2024; Rutherford et al.,, 2024; Liang et al.,, 2018; Dalton et al.,, 2020; Makoviychuk et al.,, 2021; Lange,, 2022; Lu et al.,, 2022). Most of these works rely on JAX (Bradbury et al.,, 2018; Heek et al.,, 2023; Hennigan et al.,, 2020), which enables JIT compilation, operator fusion and other components necessary for efficient vectorized code execution. These features significantly accelerate data collection by supporting rollouts in hundreds of parallelized environments. We build on these advances to scale self-supervised RL to data-rich settings.
@@ 2.3 Self-Supervised RL
Self-supervised training has enabled key breakthroughs in language modeling and computer vision (Sermanet et al.,, 2017; Zhu et al.,, 2020; Devlin et al.,, 2019; He et al.,, 2022; Mikolov et al.,, 2013). In the context of RL, by the term "self-supervised", we mean techniques that can be learned through interaction with an environment without a reward signal. Perhaps the most successful form of self-supervision has been in multi-agent games that can be rapidly simulated, such as Go and Chess, where self-play has enabled the creation of superhuman agents (Silver et al.,, 2016; 2017; Zha et al.,, 2021). When learning goal-reaching agents, another basic form of self-supervision is to relabel trajectories as successful demonstrations of the goal that was reached, even if it differs from the original commanded goal (Kaelbling,, 1993; Venkattaramanujam et al.,, 2020). This technique has seen recent adoption as “hindsight experience replay” for various deep RL algorithms (Andrychowicz et al.,, 2017; Ghosh et al.,, 2021; Chebotar et al.,, 2021; Rauber et al.,, 2021; Eysenbach et al.,, 2022).
Another perspective on self-supervision is intrinsic motivation, broadly defined as when an agent computes its own reward signal (Barto,, 2013). Intrinsic motivation methods include curiosity (Barto,, 2013; Bellemare et al.,, 2016; Baumli et al.,, 2021), surprise minimization (Berseth et al.,, 2019; Rhinehart et al.,, 2021), and empowerment (Klyubin et al.,, 2005; de Abril and Kanai,, 2018; Choi et al.,, 2021; Myers et al., 2024a, ). Closely related are skill discovery methods, which aim to construct intrinsic rewards for diverse collections of behavior (Gregor et al.,, 2016; Eysenbach et al.,, 2019; Sharma et al.,, 2020; Kim et al.,, 2021; Park et al.,, 2021; Laskin et al.,, 2022; Park et al.,, 2024). Self-supervised RL methods have been difficult to scale due to the need for many environment interactions (Franke et al.,, 2021; Mnih et al.,, 2015). This work addresses that challenge by offering a fast and scalable contrastive RL algorithm on a benchmark of diverse tasks.
@@ 2.4 RL Benchmarks
The RL community has recently started to pay greater attention to how RL research is conducted, reported, and evaluated (Henderson et al.,, 2018; Jordan et al.,, 2020; Agarwal et al.,, 2022). A key issue is a lack of reliable and efficient benchmarks: it becomes hard to rigorously compare novel methods when the number of trials needed to see statistically significant results across diverse settings ranges in the thousands of training hours (Jordan et al.,, 2024). Some benchmarks that have seen adoption include OpenAI gym/Gymnasium (Brockman et al.,, 2022; Towers et al.,, 2024), DeepMind Control Suite (Tassa et al.,, 2018), and D4RL (Fu et al.,, 2021). More recently, hardware-accelerated versions of some of these benchmarks have been proposed (Gu et al.,, 2021; Matthews et al.,, 2024; Koyamada et al.,, 2023; Makoviychuk et al.,, 2021; Nikulin et al.,, 2023). However, the RL community still lacks advanced benchmarks for goal-conditioned methods. We address this gap with JaxGCRL, significantly lowering the GCRL evaluation cost and thereby enabling impactful RL research.
@@ 3 Preliminaries
In the goal-conditioned reinforcement learning setting, an agent interacts with a controlled Markov process (CMP) ${\mathcal{M}}=({\mathcal{S}},{\mathcal{A}},p,p_{0},\gamma)$ to reach arbitrary goals (Kaelbling,, 1993; Andrychowicz et al.,, 2017; Blier et al.,, 2021). At any time $t$ the agent will observe a state $s_{t}$ and select a corresponding action $a_{t}\in{\mathcal{A}}$. The dynamics of this interaction are defined by the distribution $p(s_{t+1}\mid s_{t},a_{t})$, with an initial distribution $p_{0}(s_{0})$ over the state at the start of a trajectory, for $s_{t}\in{\mathcal{S}}$ and $a_{t}\in{\mathcal{A}}$.
For any goal $g\in{\mathcal{S}}$, we cast optimal goal-reaching as a problem of inference (Borsa et al.,, 2019; Barreto et al.,, 2022; Blier et al.,, 2021; Eysenbach et al.,, 2022): given the current state and desired goal, what is the most likely action that will bring us toward that goal? As we will see, this is equivalent to solving the Markov decision process (MDP) $\mathcal{M}_{g}$ obtained by augmenting $\mathcal{M}$ with the goal-conditioned reward $r_{g}(s_{t},a_{t})\triangleq(1-\gamma)\gamma p(s_{t+1}=g\mid s_{t},a_{t})$. Formally, a goal-conditioned policy $\pi(a\mid s,g)$ receives both the current observation of the environment as well as a goal $g\in{\mathcal{S}}$.
We denote the $k$-step action-conditioned policy distribution $p_{k}^{\pi}(s_{k}\mid s_{0},a_{0})$ as the distribution of states $k$ steps in the future given the initial state $s_{0}$ and action $a_{0}$ under $\pi$. We define the discounted state visitation distribution as $p^{\pi}_{\gamma}(s^{+}\mid s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p^{\pi}_{t}(s^{+}\mid s,a)$, which we interpret as the distribution of the state $T$ steps in the future for $T\sim\operatorname{Geom}(1-\gamma)$. This last expression is precisely the $Q$-function of the policy $\pi(\cdot\mid\cdot,g)$ for the reward $r_{g}$: $Q^{\pi}_{g}(s,a)\triangleq p^{\pi}_{\gamma}(g\mid s,a)$ (see Section D.1). For a given distribution over goals $g\sim p_{{\mathcal{G}}}$, we can now write the overall objective as
@@ 3.1 Contrastive Critic Learning
CRL is an actor-critic method that aims to solve Eq. 1. The critic is represented as a state-action-goal value function $f(s,a,g)$, which provides the likelihood of future states and how various actions influence the likelihood of future states. This function satisfies:
@@
Therefore, we can treat $f(s,a,g)$ as an approximation of the Q-function and use it to train the actor. This approach builds on previous research that frames learning of this critic as a classification problem (Eysenbach et al.,, 2022; 2021; Zheng et al.,, 2023; 2024; Farebrother et al.,, 2024; Myers et al., 2024b, ). Training is performed on batches of $(s,a,g)$ to classify whether or not $g$ is the future state corresponding to the trajectory starting from $(s,a)$. Thus, in each sample from batch $(s_{i},a_{i},g_{i})\in\mathcal{B}$ the goal $g_{i}$ is sampled from the future of the trajectory containing $(s_{i},a_{i})$.
The family of CRL algorithms consists of the following components: (a) the state-action pair and goal state representations, $\phi(s,a)$ and $\psi(g)$, respectively; (b) the critic, which is defined as an energy function $f_{\phi,\psi}(s,a,g)$, measuring some form of similarity between $\phi(s,a)$ and $\psi(g)$; and (c) a contrastive loss function, which is a function of the matrix containing the critic values $\{f_{\phi,\psi}(s_{i},a_{i},g_{j})_{i,j}\}$ over the elements of the batch $\mathcal{B}$. The base contrastive loss we study will be the infoNCE objective (Sohn,, 2016), modified to use a symmetrized (Radford et al.,, 2021) critic parameterized with $\ell_{2}$-distances (Eysenbach et al.,, 2024). The final objective for the critic can thus be expressed as:
@@
This loss contrasts each positive sample with the batch of negative samples. Other losses, which are tested in Section 5.3, are further discussed in Section A.2.

### v4

@@ Abstract
Self-supervision has the potential to transform reinforcement learning (RL), paralleling the breakthroughs it has enabled in other areas of machine learning. While self-supervised learning in other domains aims to find patterns in a fixed dataset, self-supervised goal-conditioneds reinforcement learning (GCRL) agents discover new behaviors by learning from the goals achieved during unstructured interaction with the environment. However, these methods have failed to see similar success, both due to a lack of data from slow environment simulations as well as a lack of stable algorithms. We take a step toward addressing both of these issues by releasing a high-performance codebase and benchmark (JaxGCRL) for self-supervised GCRL, enabling researchers to train agents for millions of environment steps in minutes on a single GPU. By utilizing GPU-accelerated replay buffers, environments, and a stable contrastive RL algorithm, we reduce training time by up to $22\times$. Additionally, we assess key design choices in contrastive RL, identifying those that most effectively stabilize and enhance training performance. With this approach, we provide a foundation for future research in self-supervised GCRL, enabling researchers to quickly iterate on new ideas and evaluate them in diverse and challenging environments. Code: https://github.com/MichalBortkiewicz/JaxGCRL.
@@ 1 Introduction
Self-supervised learning has significantly influenced machine learning over the last decade, transforming how research is done in domains such as natural language processing and computer vision (chen2020simple; dosovitskiy2020image; vaswani2017attention). In the context of reinforcement learning (RL), most self-supervised prior methods apply the same recipe that has been successful in other domains: learning representations or models from a large, fixed dataset (hoffmann2022training; sardana2024chinchillaoptimal; muennighoff2023scaling). However, the RL setting also enables a fundamentally different type of self-supervised learning: rather than learning from a fixed dataset (as done in NLP and computer vision), a self-supervised reinforcement learner can collect its own dataset. Thus, rather than learning a representation of a dataset, the self-supervised reinforcement learner acquires a representation of an environment or of behaviors and optimal policies therein. Self-supervised reinforcement learners may address many of the challenges that stymie today’s foundation’s models: reasoning about the consequences of actions (rajani2019explain; kwon2023grounded) (i.e., counterfactuals (bhargava2022commonsense; jin2023cladder)) and long horizon planning (bhargava2022commonsense; du2023what; guan2023leveraging).
@@
In this paper, we study self-supervised RL in an online setting: an agent interacts in an environment without a reward to learn representations, which are later used to quickly solve downstream tasks. We focus on goal-conditioned reinforcement learning (GCRL) algorithms, which aim to use these unsupervised interactions to learn policies for achieving various goals – an essential capability for multipurpose robots. Prior work has proposed several algorithms for self-supervised RL (eysenbach2022contrastive; zheng2024stabilizing; myers2023goal; myers2024learninga), including algorithms that focus on learning goals (erraqabiTemporalAbstractionsAugmentedTemporally2022). However, these methods were limited by small datasets and infrequent online interactions with the environment, which has prevented them from exploring the potential emergent properties of self-supervised reinforcement learning on large-scale data.
[…]
@@ 2.1 Goal-Conditioned Reinforcement Learning
GCRL is a special case of the general multi-task RL setting, in which the potential tasks are defined by goal states that the agent is trying to reach in an environment (kaelbling1993learning; ghosh2019learning; nair2018visual). Achieving any goal is appealing for generalist agents, as it allows for diverse behaviors without needing specific reward functions for each task (each state defines a task when seen as a goal) (schaul2015universal). GCRL techniques have seen success in domains such as robotic manipulation (andrychowicz2017hindsight; eysenbach2021clearning; ghosh2021learning; ding2022generalizing; walke2023bridgedata) and navigation (shah2023vint; levine2023learning; manderson2020visionbased; hoang2021successor). Recent work has shown that representations of goals can be imbued with additional structure to enable capabilities such as language grounding (myers2023goal; ma2023liva), compositionality (liu2023metric; myers2024learning; wang2023optimal), and planning (park2023hiql; eysenbach2024inference). We show how GCRL techniques based on contrastive learning (eysenbach2022contrastive) can be scaled with GPU acceleration to enable fast and stable training.
@@ 2.2 Accelerating Deep Reinforcement Learning
Deep RL has only recently become practical for many tasks, in part due to improvements in hardware support for these algorithms. Distributed training has enabled RL algorithms to scale across hundreds of GPUs (mnih2016asynchronous; espeholt2018impala; espeholt2020seed; hoffman2022acme). To resolve the bottleneck of environment interaction with CPU-bound environments, various GPU-accelerated environments have been proposed (matthews2024craftax; freeman2021brax; bonnet2024jumanji; rutherford2024jaxmarl; liang2018gpuaccelerated; dalton2020accelerating; makoviychuk2021isaaca; gymnax2022github; lu2022discovered). Most of these works rely on JAX (bradbury2018jax; heek2023flax; haiku2020github), which enables JIT compilation, operator fusion and other components necessary for efficient vectorized code execution. These features significantly accelerate data collection by supporting rollouts in hundreds of parallelized environments. We build on these advances to scale self-supervised RL to data-rich settings.
@@ 2.3 Self-Supervised RL
Self-supervised training has enabled key breakthroughs in language modeling and computer vision (sermanet2017timecontrastive; zhu2020s3vae; devlin2019berta; he2022masked; mikolov2013distributed). In the context of RL, by the term "self-supervised", we mean techniques that can be learned through interaction with an environment without a reward signal. Perhaps the most successful form of self-supervision has been in multi-agent games that can be rapidly simulated, such as Go and Chess, where self-play has enabled the creation of superhuman agents (silver2016mastering; silver2017mastering; zha2021douzero). When learning goal-reaching agents, another basic form of self-supervision is to relabel trajectories as successful demonstrations of the goal that was reached, even if it differs from the original commanded goal (kaelbling1993learning; venkattaramanujam2020selfsupervised). This technique has seen recent adoption as “hindsight experience replay” for various deep RL algorithms (andrychowicz2017hindsight; ghosh2021learning; chebotar2021actionable; rauber2021hindsight; eysenbach2022contrastive).
Another perspective on self-supervision is intrinsic motivation, broadly defined as when an agent computes its own reward signal (barto2013intrinsic). Intrinsic motivation methods include curiosity (barto2013intrinsic; bellemare2016unifying; baumli2021relative), surprise minimization (berseth2019smirl; rhinehart2021information), and empowerment (klyubin2005empowerment; deabril2018unified; choi2021variational; myers2024learninga). Closely related are skill discovery methods, which aim to construct intrinsic rewards for diverse collections of behavior (gregor2016variationala; eysenbach2019diversity; sharma2020dynamicsaware; kim2021unsupervised; parkLipschitzconstrainedUnsupervisedSkill2021; laskin2022cic; park2024metra). Self-supervised RL methods have been difficult to scale due to the need for many environment interactions (franke2021sampleefficient; mnih2015humanlevel). This work addresses that challenge by offering a fast and scalable contrastive RL algorithm on a benchmark of diverse tasks.
@@ 2.4 RL Benchmarks
The RL community has recently started to pay greater attention to how RL research is conducted, reported, and evaluated (henderson2018deep; jordan2020evaluating; agarwal2022deep). A key issue is a lack of reliable and efficient benchmarks: it becomes hard to rigorously compare novel methods when the number of trials needed to see statistically significant results across diverse settings ranges in the thousands of training hours (jordan2024position). Some benchmarks that have seen adoption include OpenAI gym/Gymnasium (brockman2022openai; towers2024gymnasium), DeepMind Control Suite (tassa2018deepmind), and D4RL (fu2021d4rl). More recently, hardware-accelerated versions of some of these benchmarks have been proposed (gu2021braxlines; matthews2024craftax; koyamada2023pgx; makoviychuk2021isaaca; nikulin2023xlandminigrid). However, the RL community still lacks advanced benchmarks for goal-conditioned methods. We address this gap with JaxGCRL, significantly lowering the GCRL evaluation cost and thereby enabling impactful RL research.
@@ 3 Preliminaries
In the goal-conditioned reinforcement learning setting, an agent interacts with a controlled Markov process (CMP) ${\mathcal{M}}=({\mathcal{S}},{\mathcal{A}},p,p_{0},\gamma)$ to reach arbitrary goals (kaelbling1993learning; andrychowicz2017hindsight; blier2021learning). At any time $t$ the agent will observe a state $s_{t}$ and select a corresponding action $a_{t}\in{\mathcal{A}}$. The dynamics of this interaction are defined by the distribution $p(s_{t+1}\mid s_{t},a_{t})$, with an initial distribution $p_{0}(s_{0})$ over the state at the start of a trajectory, for $s_{t}\in{\mathcal{S}}$ and $a_{t}\in{\mathcal{A}}$.
For any goal $g\in{\mathcal{S}}$, we cast optimal goal-reaching as a problem of inference (borsa2019universal; barreto2022successor; blier2021learning; eysenbach2022contrastive): given the current state and desired goal, what is the most likely action that will bring us toward that goal? As we will see, this is equivalent to solving the Markov decision process (MDP) $\mathcal{M}_{g}$ obtained by augmenting $\mathcal{M}$ with the goal-conditioned reward $r_{g}(s_{t},a_{t})\triangleq(1-\gamma)\gamma p(s_{t+1}=g\mid s_{t},a_{t})$. Formally, a goal-conditioned policy $\pi(a\mid s,g)$ receives both the current observation of the environment as well as a goal $g\in{\mathcal{S}}$.
We denote the $k$-step action-conditioned policy distribution $p_{k}^{\pi}(s_{k}\mid s_{0},a_{0})$ as the distribution of states $k$ steps in the future given the initial state $s_{0}$ and action $a_{0}$ under $\pi$. We define the discounted state visitation distribution as $p^{\pi}_{\gamma}(s^{+}\mid s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p^{\pi}_{t}(s^{+}\mid s,a)$, which we interpret as the distribution of the state $T$ steps in the future for $T\sim\operatorname{Geom}(1-\gamma)$. This last expression is precisely the $Q$-function of the policy $\pi(\cdot\mid\cdot,g)$ for the reward $r_{g}$: $Q^{\pi}_{g}(s,a)\triangleq p^{\pi}_{\gamma}(g\mid s,a)$ (see Section˜D.1). For a given distribution over goals $g\sim p_{{\mathcal{G}}}$, we can now write the overall objective as
@@ 3.1 Contrastive Critic Learning
CRL is an actor-critic method that aims to solve Eq.˜1. The critic is represented as a state-action-goal value function $f(s,a,g)$, which provides the likelihood of future states and how various actions influence the likelihood of future states. This function satisfies:
@@
Therefore, we can treat $f(s,a,g)$ as an approximation of the Q-function and use it to train the actor. This approach builds on previous research that frames learning of this critic as a classification problem (eysenbach2022contrastive; eysenbach2021clearning; zheng2023contrastive; zheng2024stabilizing; farebrother2024stop; myers2024learning). Training is performed on batches of $(s,a,g)$ to classify whether or not $g$ is the future state corresponding to the trajectory starting from $(s,a)$. Thus, in each sample from batch $(s_{i},a_{i},g_{i})\in\mathcal{B}$ the goal $g_{i}$ is sampled from the future of the trajectory containing $(s_{i},a_{i})$.
The family of CRL algorithms consists of the following components: (a) the state-action pair and goal state representations, $\phi(s,a)$ and $\psi(g)$, respectively; (b) the critic, which is defined as an energy function $f_{\phi,\psi}(s,a,g)$, measuring some form of similarity between $\phi(s,a)$ and $\psi(g)$; and (c) a contrastive loss function, which is a function of the matrix containing the critic values $\{f_{\phi,\psi}(s_{i},a_{i},g_{j})_{i,j}\}$ over the elements of the batch $\mathcal{B}$. The base contrastive loss we study will be the infoNCE objective (sohn2016improved), modified to use a symmetrized (radford2021learning) critic parameterized with $\ell_{2}$-distances (eysenbach2024inference). The final objective for the critic can thus be expressed as:
@@
This loss contrasts each positive sample with the batch of negative samples. Other losses, which are tested in Section˜5.3, are further discussed in Section˜A.2.